OpenStudy (anonymous):

-x + 3y = 12 if when y = 3

OpenStudy (anonymous):

so plug that 3 in for y and solve for x

OpenStudy (anonymous):

-x+3y=12 -x+3(3)=12 -x+9=12 -x=3 x=-3 so your answer is x=-3